### v1

Abstract
Electric vehicle charging demand prediction is important for vacant charging pile recommendation and charging infrastructure planning, thus facilitating vehicle electrification and green energy development. The performance of previous spatio-temporal studies is still far from satisfactory because the traditional graphs are difficult to model non-pairwise spatial relationships and multivariate temporal features are not adequately taken into account. To tackle these issues, we propose an attention-based heterogeneous multivariate data fusion approach (AHMDF) for citywide electric vehicle charging demand prediction, which incorporates geo-based clustered hypergraph and multivariate gated Transformer to considers both static and dynamic influences. To learn non-pairwise relationships, we cluster service areas by the types and numbers of points of interest in the areas and develop attentive hypergraph networks accordingly. Graph attention mechanisms are used for information propagation between neighboring areas. Additionally, we improve the Transformer encoder utilizing gated mechanisms so that it can selectively learn dynamic auxiliary information and temporal features. Experiments on an electric vehicle charging benchmark dataset demonstrate the effectiveness of our proposed approach compared with a broad range of competing baselines. Furthermore, we demonstrate the impact of dynamic influences on prediction results in different areas of the city and the effectiveness of our clustering method.
keywords:
[…]
1 Introduction
As the world accelerates its transition to low-carbon and environmental friendly mobility, electric vehicles (EVs) will play an important role in the future automotive market [1]. The International Energy Agency expects that policies in major economies around the world will increase the market share of electric vehicles to 50% by 2030. However, the lack of available and reasonably located charging facilities is a major constraint to the development of electric vehicles [2, 3]. The increasing development of smartphones, in-vehicle devices and information distribution systems makes governments and mangers glad to turn to accurate EV charging demand prediction for help. By predicting the demand for services in the coming future, administrators can provide users with suggestive information, such as timing and route planning, to avoid customer losses. For governments, accurate charging demand prediction, especially at the citywide level, is a groundwork for a variety of smart strategies such as charging infrastructure planning and smart power supply [4, 5, 6].
Researchers have made many efforts to enable highly accurate EV charging demand prediction. First, due to the simplicity and computational speed of the statistical method, it has been widely used in early studies to regress time-series charging demand data, but the linear assumption limits its accuracy [7, 8]. With the development of computer technology, machine learning has become an effective method for charging demand prediction, especially spatio-temporal prediction, as it can accurately learn data patterns and fuse multidimensional features. Many models introduce Recurrent neural networks to extract temporal features [9, 10] and Graph neural networks to extract spatial features [11], respectively. However, the ambiguity in the weight assignment of convolutional and recurrent operations limits their capabilities. In the latest researches, time-series attention networks (e.g., Transformer) [12] and graph attention networks [13] that are able to better determine the weights on the network have become popular tools for improving the spatio-temporal prediction accuracy. However, classical attention mechanisms focus on the entire sequence of input hidden states, whereas not all hidden states can provide enough useful information for prediction, especially in long time series. Moreover, some typical models lack consideration of external influences and are difficult to cope with complex dynamics, making prediction lagging and less accurate [14].
Figure 1: The impact of different factors on charging demand.
EV charging demand is typically influenced by many factors that can be categorized into two main groups, namely static and dynamic, as shown in Fig. 1. Static influences include neighborhood charging services and land use around the charging stations. These influences are heterogeneous to the time-series charging demands and are usually not directly acceptable as inputs to the network. For example, improvements of the road network and the smart application on mobile devices make it possible for demand in neighboring areas to influence each other. Previous studies employ graph neural networks to fuse information from adjacent nodes and achieve good results [15, 16]. Moreover, some areas have similar land uses and residents may have typical and similar charging habits within those areas [17]. For example, charging demand in commercial areas is relatively high on weekends and after work on weekdays, while it’s the opposite in office areas. Clustering similar nodes and then training the network in groups is a common method used in previous studies [18]. However, it is still a challenge to accurately cluster nodes with similar properties. And separating the network into multiple groups for training makes it difficult to share information between groups and consumes more computational resources and storage space. Dynamic influences and demand data are usually isomorphic, both characterized by changes over time. These dynamics include charging price, charging environment, and other variables that affect charging demand in real time. For example, price influences the charging choices of price-sensitive users, and temperature largely affects the power consumption and charging efficiency of electric vehicles [19]. Previous research mainly used convolution to fuse dynamic external data. Limited by the ambiguity of convolution, these external data bring a certain but weak improvement in prediction accuracy [20, 21, 22]. A more precise method for determining the significance of dynamic influences is needed.
In summary, to build an accurate citywide EV charging demand prediction model that can take into account heterogeneous multi-source influences, several challenges are emerging, which include:
  1. [1)]
  2. 1.
    Static heterogeneous data fusion: Since areas with similar attributes may have similar data variations, mining regional attribute information becomes a promising topic for improving prediction accuracy. A suitable method for clustering areas and effective networks to extract the same patterns of areas with similar land use remain to be explored.
  3. 2.
    Dynamic isomorphic data fusion: Time-series influencing data are fused with demand data mainly in a convolutional approach in previous studies, which results in computational ambiguities and uncertainties in the weights. Different variables have different contribution to prediction and a network that effectively selects and considers variables should be employed to enhance information identification.
  4. 3.
    Structure optimization of prediction network: It remains to be explored how classical neural networks can be improved to more suitably adapt to citywide EV charging demand prediction. For example, possibly not every timestamp contributes sufficiently to the predicted results, and we need to explore sparse and rational allocation of attention in temporal attention networks.
To address the above challenges, we propose an attention-based heterogeneous multivariate data fusion approach for citywide EV charging demand prediction, named AHMDF. It enables 1) classifying areas based on an adaptive clustering method and simultaneously learning similar patterns from areas with similar properties through attentive hypergraph; 2) information propagation between neighboring areas through graph attention; and 3) fusing time-series external variables and efficiently capturing temporal features through multivariate gated Transformer. We demonstrate the state-of-the-art performance of the proposed AHMDF method through a dataset collected from Shenzhen, a first-tier city in China. We also demonstrate the contribution of multiple spatio-temporal influences to accurate EV charging demand prediction through the case study.
The main contributions of this paper are as follows.
  • We propose a novel deep learning framework for citywide electric vehicle charging prediction that can comprehensively capture the spatio-temporal dependencies and effectively consider charging-related dynamic and static influences. We incorporate an area attributes fusion module, an adjacency graph fusion module, and a temporal feature fusion module to achieve accurate prediction.
  • We propose a geo-based clustered hypergraph network for considering urban properties around charging service facilities. We propose an area clustering scheme that combines Term Frequency-Inverse Document Frequency (TF-IDF) and K-means, which can adaptively determine the importance of POI categorization and then cluster urban areas by functions.
  • We propose multivariate gated Transformer to enable the integration of a variable selection network and an improved Transformer encoder. It computes the attention scores of the variables through gated residual networks (GRNs) and efficiently combines external dynamic data. Furthermore, it utilizes GRNs and Gumbel-Softmax to improve the Transformer encoder, which determines sparse attention for the timestamps and avoids the fuzzy perception of long-term attention.
The remaining article is structured as follows. Section 2 reviews the relevant literature. Section 3 presents some preliminary concepts and the problem formulation. Section 4 describes our proposed method, AHMDF, in detail. The method is validated in Section 5 with a representative open access dataset. Section 6 draws conclusions and future directions.
2 Literature review
Citywide EV charging demand prediction can be considered as a spatio-temporal regression problem, which plays an important role in urban planning, charging guidance, and grid stabilization. Early studies mainly employed statistical methods such as Linear regression, Autoregressive Integrated Moving Average (ARIMA)[23] and Least absolute shrinkage and selection operator (Lasso)[7]. Several studies have recognized that EV charging demand is associated with multi-source factors (e.g., price) and attempted to model the relationship using linear models [24]. These models are simple and computationally fast, and make important attempts at modeling time-series EV charging demand and its correlates, but are limited by their linear assumptions, and the predictions are not precise enough. The development of machine learning has quickly made it a popular method for data regression. Various methods such as Random forest and Support vector regression have been used for electric vehicle charging demand prediction and achieved good results [25, 26]. However, these simple machine learning models still on one hand lack the consideration of complex nonlinear multivariate relationships, and on the other hand it is difficult to incorporate heterogeneous data (e.g., land attributes and spatial relationships) into the model.
Neural networks and deep learning can not only accurately model temporal relationships, but also learn multivariate and multi-node features, which become a research hotspot for citywide electric vehicle charging demand prediction. Recurrent neural networks (including Long Short Term Memory (LSTM)[27, 28] and Gated Recurrent Unit (GRU)[29]) make an important contribution to the accuracy of regression. A study proposes a new hybrid LSTM neural network that contains both historical charging state sequences and time-related features, achieving excellent time-series EV charging demand prediction [9]. Graph Convolutional Neural Network (GCN) is an application of Convolutional Neural Networks (CNN) on graphs, which enables information from neighbors to be learned [11]. The combination of graph neural networks and recurrent neural networks provides the model with powerful spatio-temporal modeling capabilities [15]. A study combining GCN and GRU achieves excellent results on traffic prediction tasks [30]. Moreover, a study improved the spatio-temporal learning capability of the model by replacing the linear operation in GRU with graph convolution [31]. Attention mechanism enables the weights in the model to be determined more rationally and accurately. Transformer can not only determine the importance of temporal feature, but also parallelize the hidden state processing to achieve a faster computation speed than recurrent neural networks [12, 32]. Graph Attention Network (GAT) is a combination of GCN and attention mechanism that allow weights to be determined on graphs [13]. An air traffic density prediction study combines GAT and LSTM to make accurate and robust predictions that outperform baseline models [33]. However, recent research has shown that not every feature contributes sufficiently to the output. Classical attention mechanisms may allocate attention to features that do not contribute enough, thus limiting prediction accuracy. Gated attention network for sequence data is thus proposed [34].
Apart from EV charging demand data, more and more studies are beginning to focus on other influencing factors for assisting in accurate prediction [35, 36, 37]. One of the major influencing factors is the price, which affects the users’ choices. Incorporating price information into the model can dynamically capture user behavior and reduce prediction lag. A deep learning charging demand prediction study utilizes attention mechanisms and bi-directional recurrent neural networks to take into account the effect of price and achieve accurate predictions [38]. Our previous studies use mixed pseudo-sample meta-learning and physics-informed neural networks to model the relationship between charging demand and price, respectively, but these studies still lack consideration of other influencing factors [20, 21]. Moreover, land utilization information (e.g., points of interest) can help determine the amenities surrounding a charging station. A demand prediction study first learns spatio-temporal information by using GCN and GRU, then categorizes EV charging stations based on points of interest, and handles each class of charging stations separately using fully connected networks [18]. However, accurate methods for clustering and fusing intragroup information are still lacking. Day and night show the same cyclical variation as temperature (i.e., high temperatures during the day and low temperatures at night), and EV power consumption is closely related to the environment (e.g. EVs consume more electricity when temperatures are low). In previous studies, environment variables are used to assist in predicting traffic behavior such as parking [35], but there is still a lack of charging demand prediction studies that take environment patterns into account.
Term Frequency-Inverse Document Frequency (TF-IDF) is first used for information retrieval to obtain accurate results in areas such as searching and literature classification [39]. It determines the value of each word and identifies the keywords that are most important to the text. A study presents a novel probabilistic retrieval model, which forms a basis to interpret the TF-IDF term weights as making relevance decisions [40]. This study demonstrates the important role of TF-IDF in searching for critical information. A recent multi-view joint graph learning study applies TF-IDF to identify regions’ intrinsic key attributes and achieves outstanding results [41].
Existing graph neural network frameworks are deployed based on simple graph structures, which limits their application when dealing with complex associations of grouped data. Recently hypergraph-based methods have been proposed to solve this problem [42]. A general high-order multi-modal/multi-type modeling framework called HGNN+ is proposed to simultaneously handle multi-hop adjacencies, attribute groupings, and other hyperedge relations under a single hypergraph-based framework [43]. Hypergraph convolutional neural networks that allow information to propagate over hypergraphs have a wide range of applications in both regression and classification tasks, such as multi-task traffic prediction [44], metro passenger flow prediction [45], POI recommendation [46] and student academic performance classification [47]. Moreover, several studies utilize attention mechanisms to improve hypergraph convolutional networks to select attributes that are important for interpretation [48, 49].
In summary, deep learning enables models with powerful feature extraction capabilities, thus researchers are working on building exquisite and efficient deep learning models for accurate EV charging demand prediction. Recent researches are turning to advanced attention mechanisms to better determine the weights of the network. A growing number of studies consider multi-source factors related to EV charging demand. However, models that can comprehensively and effectively account for both static and dynamic external influences are still lacking. We still need to explore more refined approaches to learn complex spatio-temporal features for accurate citywide EV charging demand prediction.
3 Problem Formulation
First, we divide a city into sub-areas based on travel community for the convenience of urban studies. Electric vehicle charging demand is considered the same as the number of EV charging piles occupied in a given sub-area or travel community. Charging demand in the whole city is defined as $\bm{D}=\{\bm{D}_{:,1},\bm{D}_{:,2},...,\bm{D}_{:,t},...\}\in\mathbb{R}^{N_{A}\times N_{T}}$, where $N_{A}$ is the number of areas, and $N_{T}$ is the number of time steps. Demand at the t-th moment is defined as $\bm{D}_{:,t}=\{D_{1,t},D_{2,t},...,D_{N_{A},t}\}$, where $D_{N,t}$ stands for the charging demand at the $t$-th moment in the $N$-th community. Similar to the demand data, dynamic factors can be defined as $\bm{U}=\{\bm{U}_{:,1},\bm{U}_{:,2},...,\bm{U}_{:,t},...\}\in\mathbb{R}^{N_{A}\times N_{T}}$. Among the dynamic influences, we take price and temperature as examples in our experiments. It is worth noting that our methodology works for other potential dynamics as well, including traffic flow, economic indicators, etc. We slice the data using lookback mechanism with a window size $\tau$. The sliced demand data is denoted as $\bm{D}_{s}=\{\bm{D}_{:,t-\tau+1},\bm{D}_{:,t-\tau+2},...,\bm{D}_{:,t}\}$. The same operation is performed on other multi-source temporal data to get $\bm{U}_{s}$ as inputs of the model. To consider regional associations and further improve prediction accuracy, areas adjacencies (i.e., an undirected graph $\bm{G}$) and area attributes (e.g., number of different types of POIs in the area) are also considered in the model, which are heterogeneous data. Therefore, we can define the citywide electric vehicle charging demand prediction problem that considers urban region and dynamic influences as Eq. (1).
$\left\{\bm{D}_{s},\bm{U}_{s},\bm{G},\rm{POIs}\right\}\stackrel{{\scriptstyle\mathcal{F}}}{{\longrightarrow}}\bm{\hat{D}}_{t+\tau^{\prime}}$ ((1))
where $\bm{\hat{D}}_{t}$ denotes the predicted electric vehicle charging demand at the $t$-th moment.
4 Methodology
Fig. 2 illustrates the structure of AHMDF, which consists of three modules, namely a) an area attributes fusion module, in which areas with similar attributes are clustered by using TF-IDF and K-means, and grouped features are learned by using attentive hypergraph neural network; b) an adjacency graph fusion module, which allows for learning adjacency features through graph attention mechanism; and c) a temporal feature fusion module, which enables multi-source temporal variable selection and temporal feature learning through attention mechanisms and improved Transformer encoder.
Figure 2: The model structure of the proposed approach, which consists of (a). Area attributes fusion module, (b). Adjacency graph fusion module, and (c). Temporal feature fusion module.
4.1 Area Attributes Fusion Module
Similar areas usually have similar traffic characteristics, e.g., residential locations usually have relatively high demand for charging at night. The type and number of POIs can be a good indicator of the area’s utilization attributes, but it is not in the same form as demand data and cannot be used directly as an input to the model. We use POIs as the basis for area clustering to learn the similar pattern in areas groups. To enhance the accuracy of clustering, Term Frequency-Inverse Document Frequency (TF-IDF) and K-means are employed. TF-IDF is a common weighting technique used in information retrieval and mining to evaluate the importance of a word to a document. We treat a POI category as a word and an area as a document, and use this method to assess the importance of each category for the areas.
Term Frequency (TF) indicates the normalized frequency of occurrence of a term in the text, as shown in Eq. (2).
${\rm{TF}}_{i,j}=\frac{f_{i,j}}{\sum_{k}f_{k,j}}$ ((2))
where $\rm{TF}_{i,j}$ denotes the term frequency of word $i$ in document $j$, and $f_{i,j}$ denotes the number of times word $i$ appears in document $j$.
Note that some generic words do not contribute much to the theme, but rather less frequent words that convey the theme of the text, so it is not appropriate to use TF simply. The same is true for POI fusion, where some of the high-frequency POI types may not be fully representative of the area’s usage characteristics. The Inverse Document Frequency (IDF) indicates the prevalence of the word. The fewer the documents containing the word, the larger the IDF, indicating that the word has good category differentiation. The calculation of IDF is shown in Eq. (3).
${\rm IDF}_{i,j}=\log{\frac{N_{d}}{1+\left|j:i\in j\right|}}$ ((3))
where $\rm{IDF}_{i,j}$ denotes the inverse document frequency of word $i$ in document $j$, $N_{d}$ denotes the number of documents, and $\left|j:i\in j\right|$ denotes the number of documents containing word $i$.
Taken together, a high TF-IDF is expected if the frequency of a POI type is high in a specific area and the frequency of that POI is low in the entire city area. The calculation of TF-IDF is shown in Eq. (4).
${\rm{TF\text{-}IDF}}_{i,j}={\rm TF}_{i,j}\times{\rm IDF}_{i,j}$ ((4))
where $\rm{TF\text{-}IDF}_{i,j}$ denotes the Term Frequency-Inverse Document Frequency of word $i$ in document $j$.
From the above we obtain the sequence $\rm{TF\text{-}IDF}_{:,j}$ of the importance of each POI type for each area. Then, we employ K-means clustering [50] to divide the areas into a specific number of groups (denoted as $C$) based on TF-IDF.
In order to fuse features within groups, hypergraph attention mechanism is introduced. We consider all areas of the entire city as a hypergraph with incidence matrix $\bm{M}\in\left\{0,1\right\}^{\left|\nu\right|\times\left|\varepsilon\right|}$, where $\nu$ denotes the vertex set and $\varepsilon$ denotes the hyperedge set. Besides, we consider each area group as a hyperedge $\bm{e}_{i}$, which belongs to $\varepsilon$. Hyperedge $\bm{e}_{i}$ connects specific areas ${\left\{\bm{n}_{j}\mid n_{j}\in e_{i}\right\}}$. Attentive hypergraph neural network is an improvement of hypergraph convolution network, which is divided into two steps.
The first step is to propagate the information from the nodes within the cluster to the hyperedge, where we consider a entity-free hyperedge as a center node. The second step enables the node to obtain auxiliary information from the hyperedges connected to it. The calculation process is shown in Eq. (5).
$\bm{e}^{(1)}_{i}={\rm{att}}(\bm{e}_{i},\{\bm{n}_{j}\mid n_{j}\in e_{i}\})$ ((5))
$\bm{n}^{(1)}_{j}={\rm{att}}(\bm{n}_{j},\{\bm{e}^{(1)}_{i}\mid n_{j}\in e_{i}\})$
where $\bm{e}_{i}^{(1)}\in\mathbb{R}^{1\times\tau}$, $\bm{n}_{j}^{(1)}\in\mathbb{R}^{1\times\tau}$ denote high-dimensional features of hyperedges and nodes after calculation, respectively.
To avoid repetition, we take the first step as an example, showing how the update embedding $\bm{e}^{(1)}_{i}$ of a hyperedge is aggregated from itself and its connected nodes. The computational process is essentially the same as the graph attention mechanism, where features are concatenated and then feature-aware soft attention mechanism is used to determine the importance between nodes. The specific calculation process is shown in Eq. (6).
$\omega_{ij}={\rm LeakyReLU}(\bm{W}_{\omega}(\bm{e}_{i}\parallel\bm{n}_{j}))$ ((6))
${A_{ij}}={\rm Softmax}_{j}\left(\omega_{ij}\right)=\frac{{\rm exp}(\omega_{ij})}{{\textstyle\sum_{k\in\mathcal{N}_{i}}{\rm exp}(\omega_{ik})}}$
$\bm{e}^{(1)}_{i}={\rm LeakyReLU}({\textstyle\sum_{j\in\mathcal{N}_{i}}}{A_{ij}}\bm{n}_{j})$
where $\omega_{ij}$ and ${A_{ij}}$ are the relationship variable and attention value of hyperedge $\bm{e}_{i}$ and node $\bm{n}_{j}$, respectively, LeakyReLU is a nonlinear activation function, $\bm{W}_{\omega}$ is a coefficient matrix, $\parallel$" means vector concat, and $\mathcal{N}_{i}={\left\{\bm{n}_{j}\mid n_{j}\in e_{i}\right\}}$ stands for the set of i-th hyperedge’s nodes.
Through attentive hypergraph neural network, we obtain high-dimensional features $\bm{H}_{a}$ that fuse similar land use attribute areas patterns.
4.2 Adjacency Graph Fusion Module
Well-established urban road networks and information dissemination platforms allow charging demand to spill over from one area to another, so that the demand in neighboring areas is closely correlated. We employ graph attention mechanism (GAT) to model the spatial feature of urban EV charging demand. The graph attention mechanism is an improvement of graph convolution, which enables weights on the graph to be determined and information propagation to be focused. First, we define the structure of an undirected graph $\bm{G}$. We define two areas to be connected when the two areas have at least one common edge, as shown in Eq. (7).
$a_{ij}^{G}=\begin{cases}1,&\text{if area }i\text{ and area }j\text{ are spatially neighboring}\\ 0,&\text{otherwise}\end{cases}$ ((7))
where $a_{ij}^{G}$ stands for the adjacent relationship of area $i$ and area $j$.
The calculation process of GAT is shown in Eq. (8).
$\bm{n}^{(1)}_{i}={\rm{att}}(\bm{n}_{i},\{\bm{n}_{j}\mid n_{j}\sim n_{i}\})$ ((8))
where $\bm{n}^{(1)}_{i}$ is the update embedding feature of node $i$, and $n_{j}\sim n_{i}$ stands for node $i$ and node $j$ is connected.
The specific calculation steps are shown in Eq. (9), which is similar to Eq. (6). First features are concatenated, then soft attention between nodes is calculated, and finally the information is enabled to propagate on the graph based on the attention weights.
$\omega^{G}_{ij}={\rm LeakyReLU}(\bm{W}^{G}_{\omega}(\bm{n}_{i}\parallel\bm{n}_{j}))$ ((9))
${A^{G}_{ij}}={\rm Softmax}_{j}\left(\omega_{ij}\right)=\frac{{\rm exp}(\omega_{ij})}{{\textstyle\sum_{k\in\mathcal{N}_{i}}{\rm exp}(\omega_{ik})}}$
$\bm{n}^{(1)}_{i}={\rm LeakyReLU}({\textstyle\sum_{j\in\mathcal{N}_{i}}}{A_{ij}}\bm{n}_{j})$
where $\omega^{G}_{ij}$ and ${A^{G}_{ij}}$ are the relationship variable and attention value of node $\bm{n}_{i}$ and $\bm{n}_{j}$, respectively, and $\bm{W}^{G}_{\omega}$ is a coefficient matrix.
It is worth noting that using GAT to model spatial features here is actually a simplification of spatial hypergraph attention and achieves the same effect as it. When we consider a pair of nodes connected to each other as a hyperedge, the spatial feature can be treated as a hypergraph, but this brings unnecessary computation.
Through graph attention mechanism, we obtain high-dimensional features $\bm{H}_{b}$ that fuse adjacent areas patterns.
$\bm{H}_{a}$, $\bm{H}_{b}$ and sliced demand data $\bm{D}_{s}$ are added together to obtain high dimensional features $\bm{H}\in\mathbb{R}^{N_{A}\times\tau}$, as shown in Eq. (10). Note that the temporal information in them is preserved.
$\bm{H}={\rm{Add\&Norm}}(\bm{H}_{a},\bm{H}_{b},\bm{D}_{s})$ ((10))
where ${\rm{Add\&Norm}}$ stands for adding up the hidden states and layer normalization [51].
4.3 Temporal Feature Fusion Module
To accurately predict the EV charging demand, it is worthwhile to take into account external information (e.g., price, temperature) that can influence the user’s choice of charging service. Considering auxiliary information in the model is expected to effectively capture dynamic changes and reduce lags in prediction. The data structure of some external information is the same as the time-series demand data. How to incorporate this information into the model and model the temporal effects to improve prediction accuracy remains a challenge.
Figure 3: Schematic of some module structures. (a). Gated residual network, (b). Variable selection network and (c). Gated temporal attention.
Fig. 3 illustrates the structure of three building blocks of the proposed model. First, we introduce Gated Residual Network (GRN), as shown in Fig. 3(a), in order to allow the model to flexibly handle nonlinear relationships when needed. The specific calculation process of GRN is shown in Eq. (11).
$\bm{\eta}_{1}={\rm ELU}(\bm{W}_{\eta_{1}}\bm{x}+\bm{b}_{\eta_{1}})$ ((11))
$\bm{\eta}_{2}=\bm{W}_{\eta_{2}}\bm{\eta}_{1}+\bm{b}_{\eta_{2}}$
${\rm GRN}(\bm{x})={\rm{Add\&Norm}}(\bm{x},{\rm GLU}(\bm{\eta}_{2}))$
where $\bm{x}$ denotes input matrix, $\bm{W}_{\eta_{1}},\bm{W}_{\eta_{2}},\bm{b}_{\eta_{1}},\bm{b}_{\eta_{2}}$ are coefficient matrices, $\bm{\eta_{1}},\bm{\eta_{2}}$ are intermediate layers, and ELU is the Exponential Linear Unit activation function [52]. Furthermore, Gated Linear Unit (GLU) is used as gate mechanism to flexibly handle both necessary and non-necessary information for given data, shown as Eq. (12).
${\rm GLU}(\bm{\eta}_{2})=(\bm{W}_{1}^{{\rm GLU}}\bm{\eta}_{2}+\bm{b}_{1}^{{\rm GLU}})\odot\sigma(\bm{W}_{2}^{{\rm GLU}}\bm{\eta}_{2}+\bm{b}_{2}^{{\rm GLU}})$ ((12))
where $\bm{W}_{1}^{{\rm GLU}},\bm{W}_{2}^{{\rm GLU}},\bm{b}_{1}^{{\rm GLU}},\bm{b}_{2}^{{\rm GLU}}$ are coefficient matrices, $\odot$ denotes the element-wise Hadamard product, and $\sigma(.)$ stands for the sigmoid activation function. During training, dropout is applied after the second dense layer, i.e., to $\eta_{2}$ in Eq. (11).
Inspired by soft attention mechanism, we improve a variable selection network as shown in Fig. 3(b). Auxiliary information and historical demand data used for urban EV charging demand prediction are inputted. The specific contribution of a variable is determined by the network, i.e., variable selection network enables to emphasize variables that are critical to the prediction problem and ignore variables that are not. The input data $\bm{\xi}$ (including high-dimensional features $\bm{H}$, and other dynamics $\bm{U}_{s}$) are processed independently by two GRNs. A softmax function after one of the GRNs determines the variable selection weights, which are then multiplied with the embedded features. The calculation process of the variable selection network is shown in Eq. (13).
$\bm{\xi}_{1}={\rm GRN}(\bm{\xi})$ ((13))
$\bm{v}_{1}={\rm Softmax}({\rm GRN}(\bm{\xi}))$
$\bm{\xi}_{2}=\bm{v}_{1}\bm{\xi}_{1}$
where $\bm{\xi}_{1}$ is the embedded features after the first step calculation, $\bm{v}_{1}$ is the variable selection weights, and $\bm{\xi}_{2}$ denotes the output of variable selection network, which combines multiple features.
Transformer utilizes self-attention mechanism to model sequences, breaking through the limitation that recurrent neural networks cannot be computed in parallel, and becoming a research hotspot for time series tasks [12]. We adopt the framework of Transformer encoder and improve it by utilizing GRN and Gated Attention Network. Similarly to self-attention, we process the output of variable selection network (i.e., $\bm{\xi}_{2}$) independently with three GRNs to obtain the query, key, and value matrix $\bm{Q}$, $\bm{K}$, and $\bm{V}$.
The attention mechanism assigns weights to each input, and even irrelevant units are assigned small weights. This leads to two problems: 1) relevant units are assigned insufficient attention weights, and 2) units without contributions are assigned weights, leading to performance degradation [34]. Considering that in long sequences not every interval is sufficiently helpful for prediction, we develop gated temporal attention, in which we perform a Scaled Dot-Product Attention operation, but Gumbel-Softmax [53] is used as weight activation function instead of Softmax, as shown in Eq. (14) and Fig. 3(c).
${A_{T}}(\bm{Q},\bm{K},\bm{V})={\rm Gumbel}\mbox{-}{\rm Softmax}(\frac{\bm{Q}\bm{K}^{T}}{\sqrt{d_{k}}})\bm{V}$ ((14))
where $A_{T}$ is the output of the improved attention operation, and $\sqrt{d_{k}}$ acts as a scaling factor.
Gumbel-Softmax acts as binary gates and achieve sparse attention on long sequences. It provides the model with the ability to perform gradient descent which binary gates cannot. Gumbel-Softmax can be formulated as Eq. (15).
$\hat{\alpha}_{i}=\frac{{\rm exp}(({\rm log}(\alpha_{i})+\epsilon_{i})/\varpi)}{{\textstyle\sum_{j}{\rm exp}(({\rm log}(\alpha_{j})+\epsilon_{j})/\varpi)}}$ ((15))
where $\hat{\alpha}_{i}$ is the attention weight of $i$-th unit, $\alpha_{i}$ is $i$-th position of the input vector, $\epsilon_{i}$ is a random sample from Gumbel$(0,1)$, and $\varpi$ denotes the temperature coefficient, in other words a scaling factor.
Similar to Transformer encoder, the proposed data processing process is formulated as Eq. (16).
$\bm{\psi}_{1}={\rm{GTA}}(\bm{\psi})$ ((16))
$\bm{\psi}_{2}={\rm{Add\&Norm}}(\bm{\psi}_{1},\bm{\psi})$
$\bm{\psi}_{3}={\rm{GRN}}(\bm{\psi}_{2})$
$\bm{\psi}_{4}={\rm{Add\&Norm}}(\bm{\psi}_{2},\bm{\psi}_{3})$
where $\bm{\psi}$ is input matrix, which is equal to $\bm{\xi}_{2}$ in Eq. (13), $\bm{\psi}_{i}$ denotes the output of the $i$-th step, and $\rm GTA$ stands for gated temporal attention as Eq. (14). Note that $\bm{\psi}_{4}$ can be consider as an input to the next encoder layer, and the operation of encoder is repeated $N_{E}$ times.
Finally, a dense layer is used to decode the information and output the predictions.
5 Experiments
In this section, we first present the data description and the specific setup of the experiment. Then, we compare the proposed model with current state-of-the-art models to empirically illustrate the performance of AHMDF. The model generalizability is also illustrated. Furthermore,we conduct ablation experiments to show that each module and each factor can provide a contribution to improving the prediction accuracy. Finally, the impact of hyperparameters is analyzed experimentally.
5.1 Data Description
The proposed model AHMDF is validated on a benchmark dataset collected from June 19 to July 18, 2022 in Shenzhen, China. We divide the city into areas and update the occupancy rate of the counted charging piles within the area every 5 minutes. Consequently we can obtain 8640 timestamps for the month being counted. The areas division and the number of charging piles included in each area are shown in this link [^1] https://github.com/IntelligentSystemsLab/ST-EVCDP. We divide the training set, validation set and test set by 6:1:3 in chronological order, i.e., 19 June to 6 July for training, 7 July to 9 July for validation and 10 July to 18 July for test. A total of 450,175 POI data are collected from Amap [^2] https://lbs.amap.com/. POIs include a total of 14 categories such as restaurant, shopping, scenic areas, etc. Temperature data is also collected from a weather station in Shenzhen, which has a sampling interval of 30 minutes. We use linear interpolation to enable a 5-minute interval in the temperature data, making it the same as the demand data.
5.2 Experimental Setup
The hyperparameters are chosen after a detailed search experiment and we provide hyperparameter analysis in Subsection 5.6. The final experimental setup we chose is as follows. First, the lookback window size $\tau$ is set to 12, i.e. we provide the model with 60 minutes of historical information. Second, in K-means clustering, we cluster the area into 10 classes. Third, the temperature coefficient $\varpi$ is set to 1.5 to not only provide sparse attention to temporal information, but also to avoid focusing attention on one-hot vector. Fourth, the number of encoder components $N_{E}$ is set to 2. Fifth, the maximum epoch numbers of training process is set to 2000, and model training will early stop if the loss of the validation set does not decrease within 50 epochs. Sixth, Mean Square Error (MSE) is used as the loss function for all comparison models. Seventh, we use Adam [54] as an optimizer with a mini-batch size of 512. Seventh, we employ four metrics, i.e., Root Mean Squared Error (RMSE), R-Square Coefficient ($\rm{R}^{2}$), Relative Absolute Error (RAE), and Mean Absolute Error (MAE) to illustrate the performance of our proposed model. Finally, all experiments in this study are accomplished on a Windows personal computer containing a NVIDIA RTX 4000 GPU. We share the code for our proposed model AHMDF in this link. [^3] https://github.com/kuanghx3/AHMDF.
5.3 Comparison Experiment
In this subsection, the proposed model AHMDF will be compared with representative models to illustrate the state-of-the-art performance. The models for comparison includes 3 sequential neural networks, 2 spatial neural networks and 5 spatio-temporal neural networks, which are described in detail as follows:
  • Long Short-Term Memory (LSTM) [27, 9]: An improved recurrent neural network that can efficiently process temporal information.
  • State Space Model (SSM) [55]: A hotspot model for efficient sequence modeling by capturing key information in sequences through hidden states.
  • Temporal Fusion Transformer (TFT) [32]: A model combining high-performance multiview prediction with interpretable temporal dynamic insights based on a Transformer architecture.
  • Graph Convolutional Network (GCN) [11]: A convolutional neural network that can act directly on graphs and utilize their structural information.
  • General Hypergraph Neural Network (HGNN) [43]: A hypergraph convolution scheme to learn generalized data representations for a variety of tasks.
  • TGCN [30]: A neural network-based traffic prediction method combining graph convolutional networks (GCN) and gated recurrent units (GRU).
  • AGGRU [31]: A graph convolutional gated recurrent network. Unlike existing deep learning models, linear operations in gated recurrent neural networks are replaced by graph convolution operations.
  • BEGAN [33]: An air traffic demand prediction network combining long short-term memory (LSTM) and graph attention mechanisms (GAT).
  • HSTGCN [18]: A heterogeneous spatio-temporal graph convolutional network combining GCN and GRU for predicting EV charging demand at different spatio-temporal resolutions. In particular, it considers POI classification in the model.
  • PIAST [21]: A spatio-temporal electric vehicle charging demand prediction model using convolutional neural network and physics-informed neural network to fuse price information, which is our previous work.
Table 1: Performance of the compared models.
Metrics (×$\times 10^{-2}$) RMSE $\rm R^{2}$
Model 15min 30min 45min 60min Average 15min 30min 45min 60min Average
LSTM 3.49 4.76 5.73 6.58 5.14 89.85 82.11 74.79 67.17 78.48
SSM 3.51 4.80 5.82 6.70 5.21 89.94 82.20 74.70 66.97 78.45
TFT 3.47 4.60 5.48 6.18 4.93 88.14 81.40 73.68 66.60 77.46
GCN 3.73 4.95 5.88 6.70 5.31 86.35 77.82 70.15 62.10 74.11
HGNN 3.24 4.60 5.58 6.45 4.97 90.11 81.62 74.27 66.29 78.07
TGCN 3.59 4.47 5.73 6.53 5.08 87.14 83.98 71.60 63.72 76.61
AGGRU 3.04 4.80 5.50 6.39 4.93 92.13 79.07 76.71 69.04 79.24
BEGAN 2.95 4.41 5.45 6.40 4.80 92.35 83.85 75.65 65.00 79.21
HSTGCN 2.88 4.35 5.38 6.27 4.72 92.39 83.92 76.60 68.86 80.44
PIAST 2.87 4.41 5.40 6.28 4.74 92.52 83.63 76.49 66.02 79.67
AHMDF 2.74 4.18 5.16 5.96 4.51 92.91 84.35 77.14 69.63 81.01
Metrics (×$\times 10^{-2}$) RAE MAE
Model 15min 30min 45min 60min Average 15min 30min 45min 60min Average
LSTM 13.20 18.35 22.62 26.76 20.23 1.83 2.55 3.14 3.71 2.81
SSM 12.37 17.74 22.20 26.32 19.66 1.72 2.46 3.08 3.65 2.73
TFT 14.17 18.77 23.01 26.39 20.59 1.97 2.60 3.19 3.66 2.86
GCN 14.97 20.14 24.37 28.33 21.95 2.08 2.79 3.38 3.93 3.04
HGNN 12.05 17.85 22.34 26.59 19.71 1.67 2.48 3.10 3.69 2.73
TGCN 14.39 16.34 23.62 27.52 20.47 2.00 2.27 3.28 3.82 2.84
AGGRU 10.35 19.47 21.08 25.43 19.08 1.44 2.70 2.92 3.53 2.65
BEGAN 10.15 16.64 21.79 26.95 18.88 1.41 2.31 3.02 3.74 2.62
HSTGCN 10.15 16.42 20.99 25.23 18.20 1.41 2.28 2.91 3.50 2.52
PIAST 9.88 16.51 22.15 26.60 18.79 1.37 2.29 2.93 3.69 2.57
AHMDF 9.86 16.13 20.52 24.41 17.73 1.37 2.24 2.85 3.39 2.46
  • Bold denotes the best results. Same below.
In Table 1, we demonstrate the performance of the above comparison models and AHMDF in four prediction intervals (i.e., 15, 30, 45, and 60 minutes). Although the model does not achieve the best for every prediction interval and every metric, the enhancement is still substantial. Compared to ten representative methods, our proposed method AHMDF improves about 9.35% in RMSE, 10.01% in RAE, and 9.89% in MAE on average, and achieves the best fit in $\rm R^{2}$. Moreover, AHMDF outperforms the ten comparison models by 16.60%, 9.09%, 7.93% and 7.94% on average in 15, 30, 45, and 60 minutes prediction tasks, respectively. AHMDF performs well on relatively long time prediction (e.g., 60min), which is a challenge in prediction problems, demonstrating the contribution of temporal modules (e.g., improved transformer encoder) and multi-source heterogeneous information to improve the accuracy of EV charging demand prediction. To visually demonstrate the enhancement of our proposed model, we show some prediction results of AHMDF and the previously proposed PIAST in an example area in Fig. 4. Differences between predictions and labels are plotted to illustrate this improvement clearly. The proposed method AHMDF has more accurate predictions than PIAST, with lower prediction lags when the data fluctuates dramatically (e.g., during the evening peak in this area). Furthermore, it can be found that AHMDF has an outstanding lead in long-term prediction task. This demonstrates that the effective introduction of external information such as the areas’ attributes and weather, as well as the improvement of the network structure can enhance the long-term prediction accuracy of EV charging demand.
Figure 4: Lineplots of the prediction results and errors for an example area in different prediction interval.
In comparison models, firstly, GCN that only considers spatial adjacencies and lacks consideration of temporal factors is the worst. Second, complex models perform better than simple ones, because complex models can model nonlinear spatio-temporal relationships accurately. This is why current research favors delicate and complex networks, even though they may require more storage space and calculation. Thirdly, the consideration of influences on EV charging demand somehow brings about improvements in model performance. HGNN and HSTGCN, which incorporate area attribute information, have relatively outstanding performance in comparison models, illustrating that area attributes can help model spatial features from high dimensions and improve prediction accuracy by leveraging similar nodes. PIAST, which incorporates dynamic price changes to effectively infer user behavior, also achieves good prediction results.
We conduct additional experiments to show that our model does not just fit on a specific dataset. We randomly drop 20% of the 247 areas to create different geographic contexts in the experiments. Further, to avoid randomness, the experiments are repeated five times and their results are averaged. On these new datasets, we selected one temporal model (i.e., LSTM) and three spatio-temporal models (i.e., BEGAN, HSTGCN, and PIAST) as comparative models to demonstrate the performance of our proposed model. The results and average results for the 4 prediction intervals are showed in Table 2. In the additional experiments, the proposed model AHMDF improves by 11.79% in RMSE, 10.62% in RAE and 11.27% in MAE on average compared to the comparison model, and achieves the best fit in $\rm R^{2}$.
Table 2: Experiment results of random node dropping.
Metrics (×$\times 10^{-2}$) RMSE $\rm R^{2}$
Model 15min 30min 45min 60min Average 15min 30min 45min 60min Average
LSTM 4.38 5.42 6.27 7.02 5.77 83.30 75.86 68.48 60.96 72.15
BEGAN 2.89 4.45 5.46 6.34 4.78 92.60 82.54 75.84 68.60 79.89
HSTGCN 3.16 4.56 5.54 6.24 4.87 90.62 82.01 74.56 68.98 79.04
PIAST 2.90 4.35 5.35 6.21 4.70 92.36 84.11 76.87 69.11 80.61
AHMDF 2.69 4.08 5.05 5.83 4.41 92.81 84.50 77.09 69.63 81.01
Metrics RAE MAE
Model 15min 30min 45min 60min Average 15min 30min 45min 60min Average
LSTM 18.44 22.80 26.64 30.27 24.54 2.52 3.12 3.64 4.14 3.36
BEGAN 10.22 17.64 22.11 26.30 19.07 1.40 2.41 3.02 3.59 2.61
HSTGCN 11.81 17.99 22.48 25.77 19.51 1.62 2.46 3.07 3.52 2.67
PIAST 10.53 16.72 21.48 25.87 18.65 1.44 2.29 2.94 3.54 2.55
AHMDF 10.16 16.27 21.03 24.75 18.05 1.38 2.21 2.85 3.36 2.45
5.4 Ablation experiment
A comprehensive ablation experiment is designed to validate the contribution of each part of our proposed model AHMDF. We remove module (a), (b), and (c), price information, temperature information, and variable selection networks, respectively. Specifically, we also replace Gumbel-Softmax with Softmax in the gate temporal attention module to demonstrate its usefulness. In Table 3, we present the results of ablation experiments, which are conducted in 4 different intervals (i.e., 15, 30, 45, and 60 min). The full model, AHMDF, improves about 10.04% in RMSE, 15.85% in RAE, and 15.85% in MAE on average. In the 60-minute prediction, the proposed model AHMDF is weakly worse than “# Gumbel-Softmax” in $\rm R^{2}$, but better in the other three metrics. We still argue that the ablation experiment can be passed because its overall performance is not bad. Different modules and information contribute differently to improving prediction accuracy. First, Module (b) (i.e. adjacency graph fusion module) contributes the most to the prediction, demonstrating the importance of learning adjacency features. Second, Module (c) (i.e. temporal feature fusion module) is helpful in improving long-time prediction accuracy, and removing Module (c) sharply deteriorates the results as the prediction interval increases. Besides, time-series information, such as prices and temperatures, both contribute to the accuracy of long-term prediction. Third, module (a) enables the model to effectively capture the dynamic changes of the areas with same land use, thus improving prediction accuracy. Fourth, the variable selection network helps to determine the weights among the temporal variables, and the gated temporal attention allows a strong focus on important temporal features, both of which provide a non-negligible boost.
Table 3: Results of ablation experiment in four prediction intervals.
Metrics (×$\times 10^{-2}$) RMSE $\rm R^{2}$
Model 15min 30min 45min 60min Average 15min 30min 45min 60min Average
# Module (c) 3.23 4.95 5.90 6.73 5.20 90.03 76.48 67.51 57.23 72.81
# Module (b) 4.81 4.95 5.68 6.41 5.46 71.11 75.55 71.43 64.09 70.54
# Module (a) 3.39 4.52 5.58 6.20 4.92 88.90 81.86 70.83 66.20 76.95
# price 3.65 4.78 5.69 6.46 5.15 85.83 78.49 71.34 63.48 74.78
# temperature 3.37 4.60 5.56 6.41 4.98 88.37 80.94 72.69 64.65 76.66
# Var. Sele. 3.17 4.70 5.53 6.39 4.95 89.44 79.42 73.36 65.15 76.84
# Gumbel-Softmax 2.80 4.19 5.19 5.98 4.54 92.40 84.28 76.59 69.67 80.73
Full 2.74 4.18 5.16 5.96 4.51 92.91 84.35 77.14 69.63 81.01
Metrics (×$\times 10^{-2}$) MAE RAE
Model 15min 30min 45min 60min Average 15min 30min 45min 60min Average
# Module (c) 12.49 20.49 25.07 29.46 21.88 1.73 2.84 3.48 4.09 3.03
# Module (b) 22.60 20.80 23.60 27.30 23.58 3.14 2.89 3.27 3.79 3.27
# Module (a) 14.16 18.47 23.67 26.38 20.67 1.96 2.56 3.28 3.66 2.87
# price 16.27 20.57 24.21 27.82 22.22 2.26 2.85 3.36 3.86 3.08
# temperature 14.20 19.05 23.16 27.33 20.93 1.97 2.64 3.21 3.79 2.90
# Var. Sele. 13.41 20.14 23.13 27.13 20.95 1.86 2.79 3.21 3.76 2.91
# Gumbel-Softmax 10.30 16.38 21.05 24.73 18.11 1.43 2.27 2.92 3.43 2.51
Full 9.86 16.13 20.52 24.41 17.73 1.37 2.24 2.85 3.39 2.46
  • Note that "#" denotes "without", and "Var. Sele." denotes Variable selection network.
  • "# Gumbel-Softmax" means replace Gumbel-Softmax with Softmax in Gated temporal attention module.
Furthermore, in Fig. 5, we visualize the performance decline in each area in long-term prediction (i.e., 60 minutes) due to the removal of dynamic factors. The removal of price and temperature has led to a decline in overall urban EV charging demand prediction performance, which confirms the correlation between dynamic factors and charging demand. In our case, price information contributes more to prediction accuracy than temperature information, which can be concluded in Table 3 and Fig. 5. In addition, different dynamic information contributes differently to the prediction accuracy in the city. Price information contributes strongly to the residential areas (e.g., western and northwestern Shenzhen), but weakly contributes, or even brings a weak negative effect to suburban areas (e.g., eastern and northern tourist areas). In contrast, temperature information makes a positive contribution to the suburban and tourist areas, as well as to other areas of the city. This is in line with common sense. Residents’ charging in their residential areas is sensitive to price, while leisure and recreational activities are affected by the weather easily. It shows that our model can effectively learn the pattern of urban residents’ travel behavior influenced by dynamic factors.
Figure 5: Heatmap of 60-min prediction performance declines by removing factors. Errors are measured in RMSE.
5.5 Spatial correlation analysis
Figure 6: (a).Heat map of Pearson’s correlation coefficient; (b). Connections in our hypergraph.
To illustrate the effectiveness of our geo-based clustering method, we show the true correlation of 247 areas’ occupancies and the connectivity relationship obtained by our clustering method in Fig. 6. Fig. 6(a) illustrates the areas that show a strong positive correlation. We consider a relatively strong correlation when the Pearson correlation coefficient between the two areas is greater than or equal to 0.4. Fig. 6(b) illustrates the relations connected by hyperedges in our hypergraph. A black square indicates that the two areas are connected by a hyperedge, and conversely white one indicates that they are not connected. The two images are similar in many places, which shows that our clustering method can effectively capture spatial data correlation. To give concrete examples, the red box shows some residential areas in Luohu District, Shenzhen, and the purple box shows some commercial areas in Futian District, Shenzhen. Areas with the same land use type have similar patterns of changing charging demand as shown in Fig. 6(a). Corresponding to that, our approach makes good use of POIs to cluster areas with similar land use attributes and construct hyperedge connectivity relationships.
5.6 Hyperparameter analysis
Figure 7: Impact of the clustering number on the results.
Figure 8: Impact of hyperparameters (i.e., $N_{E}$ and $\varpi$) on prediction accuracy.
In this subsection, we discuss the impact of value choices of number of adaptive areas clusters $C$, number of encoder components $N_{E}$ and temperature coefficient of Gumbel-Softmax $\varpi$ on the results. First, as shown in Fig. 7, despite $C$ being set differently, the area adaptive clustering roughly matches reality. The southern urban built-up area, developing north-western, north-eastern and central parts, and forests, reservoirs and city parks are clustered separately. However, when $C$ is too small (e.g., $C=6$), the method has difficulty in distinguishing between urban sub-centers and suburban tourist areas. Besides, when $C$ is too large, the method tends to cut the city into too many patches, e.g., as many as six types of functional areas in the southern CBD area when $C=14$, which is unfavorable for subsequent hypergraph studies. Second, we determine the appropriate hyperparameters by searching experiment, as shown in Fig. 8. Our experiment results demonstrate the robustness of the proposed model AHMDF, as it is insensitive to the change of hyperparameters. $N_{E}$ affects the depth of the network. Networks that are too shallow may have difficulty capturing complex nonlinear relationships, while networks that are too deep tend to cause overfitting and waste of computational resources. $\varpi$ determines the number of temporal features that Gumbel-Softmax focuses on. A too small $\varpi$ makes the model tend to one-hot attention, causing insufficient feature attention, while too large $\varpi$ causes Gumbel-Softmax to converge to softmax, causing fuzzy attention that we wish to avoid from the beginning. Therefore, we choose the hyperparameter setting that is most favorable for performance, as shown in Subsection 5.2.
In summary, we propose AHMDF, a deep learning model that incorporates dynamic and static influences, for accurate citywide EV charging demand prediction. The experimental results empirically show that AHMDF achieves state-of-the-art, with an average improvement of 9.75% in three metrics and achieves the best fit in $\rm R^{2}$. Specifically, our method performs well on relatively long time prediction. The ablation experiments demonstrate the joint contribution of each module, and the full model achieves a 12.26% improvement over the ablation models. We also explore the role of dynamic factors in improving the prediction performance of different city areas. Furthermore, we experimentally explored the impact of hyperparameters.
6 Conclusion
Accurate EV charging demand prediction is an important foundation for smart grids and transportation systems, helping to optimize the grid and guide the demand. How to incorporate multi-source heterogeneous external data into the model so that it can have a positive impact on improving the accuracy of EV charging demand prediction is a current research hotspot. To address the remaining research challenges, we propose an attention-based approach named AHMDF that incorporates attentive hypergraph neural network for classified areas groups, adjacency graph attention and an improved Transformer encoder. AHMDF (1) outperforms 10 representative methods by 9.75% on average in three metrics and achieves the best fit in $\rm R^{2}$; (2) outperforms 7 ablation models comprehensively, demonstrating the enhancement brought by each introduced multi-source data or module. Moreover, we analyze the effect of hyperparameters and determined the preferred values in our case.
Furthermore, our research provides some perspectives on the management of smart energy and electric vehicle charging. (1) Achieving accurate charging demand prediction requires not only a focus on the charging data itself, but also on the external factors that influence EV charging choices. Therefore, fusion of information from multiple sources at the government or manager level will be of significant help in collaborative and refined management. (2) Incorporating external static and dynamic influences in the model may be of great help in achieving accurate long term prediction. (3) Among the dynamic influences, price information contributes significantly to the accurate prediction of residential areas, while temperature information contributes to that of suburban and tourist areas. (4) Associating charging facilities with similar properties in the surrounding helps in accurate node information propagation and collaborative area management.
Future research directions could be in the following aspects. (1) To explore more factors related to EV charging demand and their impacts, in order to capture the dynamics comprehensively and accurately. (2) To validate the proposed approach on other datasets including different EV usage scenarios, temperatures, prices, etc. (3) To build transferable networks that allow knowledge and prediction models to be transferred between regions or cities and make accurate demand prediction.
[…]